OpenStudy (anonymous):

square root of 16a - square root of 49a + square root of 81a

OpenStudy (anonymous):

Just take the square root of the numbers inside to the radicals and leave the "a"s to get 4 root a minus 7 root a plus 9 root a. Then you just treat a as a variable like x and you get 6 root a.

OpenStudy (anonymous):

thank you so much :) @Potatocracy

OpenStudy (anonymous):

\[\sqrt{16a}-\sqrt{49a}+\sqrt{81a}\]brake the square roots\[\sqrt{4^2 \times a}-\sqrt{7^2 \times a}+\sqrt{9^2 \times a}\]\[4\sqrt{a}-7\sqrt{a}+9\sqrt{a}\]\[(4-7+9)\sqrt{a}\]\[6\sqrt{a}\]

OpenStudy (anonymous):

thank you also :) @WHAT?!

OpenStudy (anonymous):

Anytime!